iosangel:

The general form of an parabola is 2x^2−12x−3y+12=0 What is the standard form of the parabola?

iosangel:

tis is what i got 

surjithayer:

\[3y=2x^2-12x+12=2(x^2-6x)+12\] \[3y=2(x^2-6x+9-9)+12\] \[3y=2(x-3)^2-18+12\] \[3y+6=2(x-3)^2\] \[3(y+2)=2(x-3)^2\] \[y+2=\frac{2}{3}(x-3)^2\]

surjithayer:

you are correct